OpenStudy (akeem):

I think of a number, add 7 and then multiply by 4 my answer is 80. What was my number?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (mhchen):

Any ideas?

OpenStudy (will.h):

Lol

OpenStudy (will.h):

My turn Any ideas?

OpenStudy (akeem):

um x+28=80, 4(x+7)=80, 4x+7=80, 4x+28=80 Explain your answers, and then find the value of X !

OpenStudy (skullpatrol):

Which one do you think it is @akem?

OpenStudy (mhchen):

4(x+28) = 4(x) + 4(28) NOT 4(x) + (28)

OpenStudy (mhchen):

So which one do you think it is

OpenStudy (akeem):

Well @Skullpatrol i think it could be all of them i'm just not sure how to really explain it and just so you know this is High level !

OpenStudy (sshayer):

4(x+7)=80 x+7=80/4=.. x=...

OpenStudy (skullpatrol):

Start by letting the number you are thinking of be called x. Let x = the number

OpenStudy (akeem):

so then would X = 20 since 80/4=20 ?

OpenStudy (skullpatrol):

nope.

OpenStudy (akeem):

@Sshayer ?

OpenStudy (akeem):

why not @skullpatrol ?

OpenStudy (sshayer):

x+7=20 x=...

OpenStudy (skullpatrol):

What about the "add 7" part @akeem?

OpenStudy (mhchen):

Your initial equation was correct. 4(x+7)=80 (x+7) = 80/4

OpenStudy (akeem):

@Mhchen i thought of a number then i multiplied it by 4 what was my number ?

OpenStudy (mhchen):

No it was "I thought of a number, add 7, multiplied it by 4, my answer is 80"

OpenStudy (akeem):

yea mybad

OpenStudy (mhchen):

Yeah, so according to that wording, you added 7: x+7 Then you multiplied by 4 4 * (x + 7) Then that whole thing is 80 4 * (x+7) = 80 So then you can dividing BOTH sides by 4. x + 7 = 80 / 4 Then you can subtract BOTH sides by 7. x = (80/4) - 7 Do you know how to simplify it from here?

OpenStudy (akeem):

yea i think so know i just do 80/4-7 right ?

OpenStudy (akeem):

which is 13 so X=13 ?

OpenStudy (skullpatrol):

$$\huge \text{Good job!}$$

OpenStudy (akeem):

@skullpartrol thank you for helping

OpenStudy (akeem):

@mchen thank you as well

OpenStudy (skullpatrol):

Thanks for trying to learn :-)

OpenStudy (akeem):

Welcome !